OpenStudy (anonymous):

f(x)=1/2x+5, find a formula for the inverse function. is it f^-1=(y-5)*2??

OpenStudy (anonymous):

Is it? \[f(x) = \frac{1}{2}x + 5\]

OpenStudy (anonymous):

Yes you are right @hoya

OpenStudy (anonymous):

What you can do more is to replace by y with x in the final answer: \[f^{-1}(x) = 2 (x -5) = 2x - 10\]

OpenStudy (anonymous):

thank you.